OpenStudy (anonymous):

a car travelling at 22.4 m/s skids to a stop in 2.55seconds. Determine the skidding distance of the car. (assume uniform acceleration)

OpenStudy (anonymous):

Use the Equation v = u +at to get a and then v^2 - u ^2 = 2as

OpenStudy (anonymous):

what is u for?

OpenStudy (anonymous):

d=(Vo+V)*t/2 d=(22.4m/s+0m/s)*2.55s/2 d=28.56m

OpenStudy (anonymous):

in his equation, u is original velocity (22.4m/s) and v is final (0m/s). I use Vo instead of U

OpenStudy (anonymous):

tnx :)

OpenStudy (anonymous):

i dont get it.

OpenStudy (anonymous):

28.56 is the answer?

OpenStudy (anonymous):

hey